ABC is a right angkd triangk in which LA = 90° and AB

= AC. Fi.nd LB and LC.

Grade:12th pass

2 Answers

Pawan Prajapati
3 years ago
As AB = AC [Given] L B = L C ...(i) [Angles opposite to equal sides of a triangle are equal.] In ru\BC, L A + L B + L C = 180° [Sum of angles of a triangle is 180°.) goo + L B + L B = 180° 2L B = 90° L B = 45° L B = L C = 45°.
Ram Kushwah
3 years ago
In the Δ ABC AB=AC
The angles opposite the equal sides of triangle are equal so
∠C=∠B
 
let ∠C=∠B=x
Also the sum of all three angles of a triangle is 180°
 
So ∠A+∠B +∠C=180
90+x+x=180
2x=180-90=90
Thus x=45°
SO ∠B =∠C=45°